A valószínűségszámítás és matematikai statisztika 2. dolgozat (december 9.)

Feladatok

1.  2.  3.  4.  5.  6.
Agócs László Péter  −  ±  −  −  +  +
Arany Gábor  ±  −  −  −  −  +
Bálint Levente  −  −  −  −  −  −
Csipkés László  ±  −  −  −  +  +
Eszenyi László  −  ±  +  +  +  +
Ferenczik Zoltán  −  −  +  −  +  −
Fogarasi Edina  ±  ±  +  −  +  +
Gerják Balázs  ±  −  −  ±  +  −
Juhász Tibor  ±  −  +  ±  +  −
Kis Bertalan  −  +  +  ±  +  −
Kocsis Gergely  −  ±  +  +  +  +
Kovács György  ±  −  −  −  +  +
Lázár Gábor  −  −  −  −  +  +
Nagy Csaba  +  −  +  −  +  +
Paluska Mihály  ±  −  −  ±  +  +
Pári Viktor  −  −  −  −  −  −
Pelejtei Gergely  −  −  +  −  +  +
Sipos Nándor  −  −  +  −  +  +
Szakácsi János  −  −  +  −  +  −
Szalontai Gábor  +  ±  +  ±  +  +
Szalontai József  +  −  −  −  +  +
Tóth János  +  ±  +  ±  +  +
Tóth Tamás  +  −  +  ±  +  −
Virág Tamás  ±  −  +  −  +  −
Zabos Katalin  ±  ±  +  ±  +  +

Feladatok

1.  2.  3.  4.  5.  6.
Dancsi József  −  −  −  −  −  −
Emődi Szabolcs Árpád  −  −  −  −  ±  +
Fülöp Arnold  −  ±  +  −  −  −
Hamza Péter  ±  −  −  −  +  +
Hank Sándor  −  −  −  −  −  −
Lörinczi József  −  −  +  ±  +  +
Molnár Angéla  +  −  −  −  +  +
Molnár Tibor  +  −  −  −  +  +
Nagy Gábor  −  −  +  ±  +  +

Nagy Zsolt

Páli Gábor János ± − ± − + +

Paper Andrea

Papp Gyula  −  −  ±  ±  +  −
Szabó Zsolt  ±  ±  +  ±  +  −
Szevelényi Attila  ±  −  −  −  ±  −

Tilki Csaba Tordai Jácint